A certain product has supply and demand functions given by p=7q+21 and p=329-4q, respectively, where p is the pice in dollars and q is the quantity supplied or demanded at price p. What price give market equilibrium?
Answer by Boreal(15235)

Set 7q+21=329-4q, since each side is equal to the same p.
11q=308
q=28
Solve for p
p=7(28)+21
=$217.
check with other
329-4(28)=$217.